he market for vanilla ice cream is given by the following information: 𝑄 𝑑 = 800 βˆ’ 30𝑃 𝑣 + 10𝑃 𝑐 𝑄 𝑠 = 250 + 30𝑃 𝑣 βˆ’ 10π‘ƒπ‘š Where 𝑄 𝑑 is the quantity demanded, 𝑄 𝑠 is quantity supplied, 𝑃 𝑣 is the price of vanilla ice cream, 𝑃 𝑐 is the price of chocolate ice cream and π‘ƒπ‘š is the price of milk.




1
Expert's answer
2021-03-23T07:30:18-0400

In equilibrium Qd = Qs, so:

800 βˆ’ 30𝑃 𝑣 + 10𝑃 𝑐 = 250 + 30𝑃 𝑣 βˆ’ 10π‘ƒπ‘š,

60Pv = 10Pc + 10Pm + 550,

Pv = (Pc + Pm + 55)/6.